A solution is made up by adding 12.56g of sodium benzoate to 250.0mL of water. Calculate the pH of the resulting solution.
1
Expert's answer
2020-09-29T06:28:42-0400

"C_6H_5COO^-+H_2O\to C_6H_5COOH+OH^-"

"K_b=\frac{[C_6H_5COOH][OH^-]}{[C_6H_5COO^-]}=1.6\times10^{-10}"

"[C_6H_5COO^]=\frac{12.56}{144}=0.087M"

"K_b=\frac{C\times C}{0.087}=1.6\times10^{-10}"

"C=[OH^-]=0.37\times10^{-5}M"

"pOH=-log[OH^-]=5.43"

"pH=14-pOH=14-5.43=8.57"